OpenStudy (anonymous):

What sine function represents an amplitude of 4, a period of pi over 2, no horizontal shift, and a vertical shift of −3?

OpenStudy (anonymous):

f(x) = −3 sin 4x + 4 f(x) = 4 sin 4x − 3 f(x) = 4 sin pi over 2x − 3 f(x) = −3 sin pi over 2x + 4

OpenStudy (anonymous):

c

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

3 horizontal shift means option b and c selected and c has the given period